OpenStudy (idku):

Found part 1, troubled by part 2.       In 2006, NASA’s Mars Odyssey orbiter detected violent gas eruptions on Mars, where the acceleration due to gravity is 3.7 m/s2. The jets throw sand and dust about 79.0 m above the surface.

OpenStudy (idku):

What is the speed of the material just as it leaves the surface? I found that speed to be: |v| = \(\color{red}{24.161~\rm m/s}\) (this is correct)

OpenStudy (idku):

and this is where I am having an issue: (part 2) Scientists estimate that the jets originate as high-pressure gas speeds through vents just underground at about 160 km/h. How much energy per kilogram of material is lost due to nonconservative forces as the high-speed matter forces its way to the surface and into the air?

OpenStudy (idku):

Oh, so you just apply the formula for kinetic energy (0.5mv^2), with the force of this gas?

OpenStudy (idku):

But, where did you get 23.5584 ?

OpenStudy (janu16):

that was the answer for part a but your is just little different than mine

OpenStudy (idku):

So, basically ? \(KE=\frac{1}{2}m(\Delta v)^2=\frac{1}{2}m(v_f-v_o)^2\)

OpenStudy (janu16):

mhm

OpenStudy (idku):

what does that stand for?

OpenStudy (janu16):

what?

OpenStudy (idku):

mhm

OpenStudy (janu16):

are you talking about mhm? tht means yea

OpenStudy (idku):

Oh, ok.

OpenStudy (idku):

I'm not good with American slang.

OpenStudy (janu16):

lol not really a slang you know when people go like hmm so same thing

OpenStudy (idku):

Alright.

OpenStudy (idku):

I used my number for speed, 24.161 and I got 205.71 J/kg

OpenStudy (idku):

wow, that was wrong. I guess your number was more precise.

OpenStudy (janu16):

i mean idk but thats what i came up with lol

OpenStudy (idku):

218.11

OpenStudy (idku):

and that is also wrong

OpenStudy (janu16):

ohh really? what was the right number?

OpenStudy (idku):

Will see if I can try one more time to get it right.

OpenStudy (janu16):

alright. i gtg now.

OpenStudy (idku):

E (Joules) = 1/2 × (m kg) × (v m/s)^2 E (Joules) ÷ (m kg) = 1/2 × (v m/s)^2 So maybe, E (Joules) ÷ (m kg) = 1/2 × (\(v_f\) m/s)^2 - 1/2 × (\(v_i\) m/s)^2

OpenStudy (idku):

Yup! 710 J/kg

OpenStudy (idku):

it is the change in energy, not a change in velocity ... We did, E (Joules) ÷ (m kg) = 1/2 × (\(v_f-v_i\) m/s)^2 when should have done E (Joules) ÷ (m kg) = 1/2 × (\(v_f\) m/s)^2 - 1/2 × (\(v_i\) m/s)^2

OpenStudy (idku):

we find the gain -710 J/kg, so the loss is 710 (positive) J/kg